A silversmith mixed 15g of an 80% silver alloy with 50g of a silver alloy of 25%. What is the percent concentration of resulting alloy?
Answer by stanbon(75887)

A silversmith mixed 15g of an 80% silver alloy with 50g of a silver alloy of 25%. What is the percent concentration of resulting alloy?
---------------------
Equation:
silver + silver = silver
0.80*15 + 0.25*50 = x(15+50)
---------
80*15 + 25*50 = 6500x
-----
1200 + 1250 = 6500x
2450 = 6500x
x = 0.3769 = 37.69%
---------------------------
Cheers,
Stan H.
==================